Skip to main content

IODP Expedition 379

Amundsen Sea West Antarctic Ice Sheet History

Week 9 Report (10–16 March 2019)

Download PDF

The first part of this week was spent in transit to Puerto Williams, Chile, to evacuate an injured crew member. After this was completed, we transited to Punta Arenas, Chile, and started port call activities.


This week began while we were in transit to a point south of Cape Horn for a medical evacuation. As the transit progressed, the evacuation location was changed to Puerto Williams, Chile. We reached the Cape Horn pilot station at 1400 h on 12 March, two pilots boarded at 1503 h, and we proceeded to Puerto Williams. At 0006 h on 13 March, we arrived in Puerto Williams and anchored. Chilean officials (agent, immigration, port authority, police, and translator) boarded at 0020 h and departed at 0045 h. Later that morning, the helicopter for the medical evacuation landed on the ship at 0816 h and departed at 0838 h with the injured crew member. At 1009 h on 13 March, we departed Puerto Williams and spent the rest of the day in transit to Punta Arenas. We arrived at the Mardones Dock in Punta Arenas at 2000 h on 14 March and went through immigration and customs. Port call activities started on 15 March—we prepared the departing surface shipments and loaded incoming shipments (air freight as well as three containers of surface freight). On 16 March, we distributed incoming shipments, loaded fresh produce, and prepared the laboratories for the next expedition (reporting, cleaning).

Science Results

This week scientists finalized nearly all remaining reports, participated in group meetings to discuss postcruise research, and submitted revised postcruise research plans.


The Outreach Officers continued to document the expedition with photos, videos, writing, and comics. Social media was updated with comics and photographs. Completed this week were comics about the laboratories, scientists’ individual goals, and one about a scientist and her research in Marie Byrd Land. We contributed an illustration from Antarctica to the #kidlit4climate campaign. We are currently in discussions with Earther and Scientific American about use of the Expedition 379 comics. Interviews and photographs for postcruise comics about the crew have been completed. A live interactive event was conducted with the University of Perpignan, France, which also included French media; follow-up media articles and a photography exhibition are planned in France. Preparations have been completed for a live event with the IODP Germany Conference at their outreach event on March 18. Two blogs were published and promoted by the British Antarctic Survey. We continued to post blogs about scientists’ work on the expedition. Interviews and filming for BBC video and for postcruise video projects were completed. We participated in planning for AGU on 50 years of IODP outreach in cooperation with other outreach specialists.

Technical Support and HSE Activities

Staff engaged in supporting end-of-expedition science operations and wind-down activities, as well as port call support and logistics for upcoming Expedition 382.


  • Work continued for offgoing shipments.
  • Loaded and distributed Expedition 382’s surface freight, and packed excess supplies and sent them to local storage.

Laboratory Activities

  • Physical Properties:
    • NGR: Troubleshooting continued on detector #7 issues.
  • Underway Geophysics:
    • Collected magnetometer data on the transit to Puerto Williams.

Application Support Activities

  • Java 11:
    • Migrations for AuditLib, AutherLib, CEP (Common Entry Point), DBOverviewLib, DescinfoLib, DescLib, DPlib (Data Publishing), FilesLib, LoreLib, MonitorLib, and PitaLog.
    • Wrote instructions on how to convert services to Java 11.
    • Working on Java 11 for CitedBy utility for Publications.
  • Coordinating end-of-expedition processing with technical and science staff.
  • Assisted Chemistry Laboratory staff with storing microbiology PFT tracer measurements to LIMS.

IT Support Activities

  • Started packing Iridium communications equipment for return.
  • Collected all old hand scanners that are being swapped out with newer units.
  • Continued working with Extreme regarding ongoing issue with scientist laptop self-registration.
  • Additional UPS battery failure on UPS in VSAT Network Locker: both VSAT UPS systems in that space are now not operational, but system is still protected by other Ship’s Main UPS.
  • Worked with Rignet regarding ongoing issues with our Bow antenna system, which is still not tracking properly on occasion. Additional service call might be required. Opened up ticket with Rignet to adjust our power settings back to normal levels since we are now out of the Amundsen Sea.
  • Failure with Cisco Switch-01 in VSAT network locker resulted in outage for ~1 h. Repaired and resolved.

HSE Activities

  • Weekly fire and boat drill was held as scheduled.